Apply the method of separation of variables to solve the following Dirichlet boundary value problem
uxx +uyy = 0, 0 <x<a, 0<y <b,
u(x,0) = f(x), u(x,b) = 0, 0 ≤ x ≤ a,
u(0,y) = 0, u(a,y) = 0, 0 ≤ y ≤ b.
1
Expert's answer
2020-09-01T18:32:08-0400

"u_{xx}+u_{yy}=0, x\in(0,a), y\in(0,b)\\
u(0,y)=0,u_x(a,y)=0,y\in[0,b]\\
u(x,0)=f(x), u(x,b)=0, x\in[0,a]\\
u(x,y)=X(x)Y(y)\\
u_x=X'Y, u_{xx}=X''Y\\
u_y=XY',u_{yy}=XY''\\"

the separation of variables and used

"u(0,y)=0,u_x(a,y)=0,y\in[0,b]"

we have

"X''(x)+\lambda X(x)=0, 0<x<a, X(0)=X'(a)=0"

eigenvalues

"\lambda_k=(\frac{2k+1}{a}\frac{\pi}{2})^2, k\in Z_+"

eigenfunctions


"X_k(x)=\sin\frac{2k+1}{a}\frac{\pi}{2}x, k\in Z_+"

"Y" find from

"Y''_k-\lambda_kY_k=0, k\in Z_+\\
Y_k(y)=A_ke^{\sqrt{\lambda_k}y}+B_ke^{-\sqrt{\lambda_k}y}, k\in Z_+"

solution of problem is

"u(x,y)=\sum\limits_{k=0}^{\infty}(A_ke^{\frac{2k+1}{a}\frac{\pi}{2}y}+B_ke^{-\frac{2k+1}{a}\frac{\pi}{2}y})\sin\frac{2k+1}{a}\frac{\pi}{2}x"


"u(x,y)\to u(x,0)=f(x), u(x,b)=0, x\in[0,a]"

"\left\{\begin{matrix}
 
\sum\limits_{k=0}^{\infty}(A_k+B_k)\sin\frac{2k+1}{a}\frac{\pi}{2}x=f(x)\\
\sum\limits_{k=0}^{\infty}(A_ke^{\frac{2k+1}{a}\frac{\pi}{2}b}+B_ke^{-\frac{2k+1}{a}\frac{\pi}{2}b})\sin\frac{2k+1}{a}\frac{\pi}{2}x=0, \end{matrix}\right.\\
x\in[0,a]\\

\left\{\begin{matrix}
A_k+B_k=f_k\\
A_ke^{\frac{2k+1}{a}\frac{\pi}{2}b}+B_ke^{-\frac{2k+1}{a}\frac{\pi}{2}b}=0, k\in Z_+

 
\end{matrix}\right."

where

"f_k=\frac{2}{a}\int^{a}_{0}f(z)\sin\frac{2k+1}{a}\frac{\pi}{2}zdz, k\in Z_+"

Then

"A_k=\frac{-f_ke^{-\frac{2k+1}{a}\frac{\pi}{2}b}}{2sh\frac{2k+1}{a}\frac{\pi}{2}b}\\
B_k=\frac{f_ke^{-\frac{2k+1}{a}\frac{\pi}{2}b}}{2sh\frac{2k+1}{a}\frac{\pi}{2}b}, k\in Z_+"

"u(x,y)=\sum ^{\infty}_{k=0}\frac{f_ksh{\frac{2k+1}{a}\frac{\pi}{2}(b-y)}}{sh\frac{2k+1}{a}\frac{\pi}{2}y}\sin\frac{(2k+1)\pi}{2a}x, \\
x\in(0,a), y\in(0,b)"